# Large deviations for the smallest eigenvalue of a deformed GOE with an outlier

### v1

@@ 3.1. Properties of the rate function
We can now prove the approximation and convexity properties. Recall that $\eta_{\nu}$ is the transport map such that $\nu=\eta_{\nu}\#1_{[0,1]}\mathrm{d}x$.
@@
Indeed the second term is clearly positive since
@@
$-\inf_{[x,+\infty)}I_{\nu_{\varepsilon}^{-},t}^{\Lambda}\leq\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\\ \leq-\inf_{[x,+\infty)}I_{\nu_{\varepsilon}^{+},t}^{\Lambda}.$
@@
$-\inf_{(x,+\infty)}I_{\nu,t}^{\Lambda}\leq\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\\ \leq-\inf_{[x,+\infty)}I_{\nu,t}^{\Lambda}.$ (3.4)
@@
$-\inf_{(-\infty,x]}I_{\nu_{\epsilon}^{+},\Lambda}\leq\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\leq x\right)\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\leq x\right)\\ \leq-\inf_{(-\infty,x]}I_{\nu_{\varepsilon}^{-},\Lambda}.$
@@ 4.1. The case of the GOE with an outlier
$H_{0,t}:x\mapsto x+\frac{t}{x}.$
The characteristic has a shock at $x=-\sqrt{t}$ and the corresponding value of $H_{0,t}$ is $H_{0,t}(x)=-2\sqrt{t}$. The support of $\sigma_{t}$ is therefore $[-2\sqrt{t},2\sqrt{t}]$. Let $\omega_{0,t}$ be the inverse of $H_{0,t}$ on $(-\infty,-2\sqrt{t}]$, namely
@@
With the normalization of [Mai07], taking $\beta=1$ gives a LDP for the smallest eigenvalue of $G_{N}+\Lambda e_{1}e_{1}^{T}$ where $G_{N}$ is GOE of variance $t=1/2$. In the sequel we take $t=1/2$. Denote $\rho_{\nu,t}^{\Lambda}:=\Lambda+\frac{1}{2\Lambda}$. Let $J_{\Lambda}$ be the rate function found in [Mai07], given if $\Lambda\leq-\frac{1}{\sqrt{2}}$ by
@@
${M_{\Lambda}}(x):=-\frac{1}{2}\int_{x}^{\rho_{\nu,t}^{\Lambda}}\sqrt{z^{2}-2}\mathrm{d}z-\Lambda(x-\rho_{\nu,t}^{\Lambda}))+\frac{1}{4}(x^{2}-\rho_{\nu,t}^{\Lambda})$
@@
$J_{\Lambda}(x):=\begin{cases}N_{\Lambda}(x)&\text{if $x\leq\rho_{\nu,t}^{\Lambda}$}\\ \int_{x}^{-\sqrt{2}}\sqrt{z^{2}-2}\mathrm{d}z&\text{if $\rho_{\nu,t}^{\Lambda}\leq x\leq-\sqrt{2}$}\\ +\infty&\text{if $x>-\sqrt{2}$},\end{cases}$
@@
${N_{\Lambda}}(x):=-\frac{1}{2}\int_{x}^{-\sqrt{2}}\sqrt{z^{2}-2}\mathrm{d}z-\Lambda x+\frac{1}{4}x^{2}+\frac{1}{4}+\frac{1}{4}\log 2+\frac{\Lambda^{2}}{2}+\frac{1}{2}\log\Lambda.$
@@
Let $x<-\sqrt{2}$. Assume that $\Lambda\leq-\frac{1}{\sqrt{2}}$, or $\Lambda\geq-\frac{1}{\sqrt{2}}$ and $x<-\sqrt{2}$. One can note that
@@
Moreover
$\partial_{x}I_{0,1/2}^{\Lambda}(x)=(\omega_{\nu,1/2}(x)-\Lambda)=-\frac{1}{2}\sqrt{x^{2}-2}+\frac{x}{2}-\Lambda=J_{\Lambda}^{\prime}(x).$
Assume that $\Lambda\geq-\frac{1}{\sqrt{2}}$ and $x>\rho_{\nu,1/2}^{\Lambda}$. By Equation (3.2),
$\partial_{x}I_{0,1/2}^{\Lambda}(x)=\omega_{\nu,1/2}(x)-\omega_{\nu,1/2}^{*}(x)=-\sqrt{x^{2}-2}=J_{\Lambda}^{\prime}(x).$
@@ 4.2. The no outlier case
Let $\rho_{\nu,{1}}^{\Lambda}:=\ell_{\nu}+G_{\nu}(\ell_{\nu})\in\mathbb{R}\cup\{-\infty\}$. For all $x\leq\ell_{\nu,1}$, let $\gamma_{1}(x)\leq\gamma_{2}(x)$ be the two solutions of the equation $H_{\nu,1}(\gamma)=x$ on $(-\infty,\ell_{\nu})$.
First observe that for all $x<\ell_{\nu,1}$,
@@ Appendix A Reminder on the free convolution
We recall some classical results on free convolution from [Bia97] that will be useful throughout this article. Recall the definitions of $v_{t}(u)$ and $\Omega_{\mu,t}$ in (1.2) and (1.1). We have the following:
@@
    Moreover, for $\eta_{1}<\cdots,\eta_{p}$, denoting $\nu_{\alpha}=\sum\alpha_{i}delta_{\eta_{i}}$, we have for every $t>0$ and $\alpha,\alpha^{\prime}\in\Delta^{p}$,
@@
Assume first that $x\leq\ell(\mu\boxplus\sigma_{t})$. First by [GM20, Lemma 11], we have
$\omega_{\mu,t}((-\infty,\ell(\mu\boxplus\sigma_{t})])\subset(-\infty,\ell(\mu)]$
and therefore $\omega\leq\ell(\mu)$. This proves that $z_{0}\leq\ell(\mu)$. Let $s\in(0,t)$. Recall that $\mu\boxplus\sigma_{t}=(\mu\boxplus\sigma_{s})\boxplus\sigma_{t-s}$. We have
@@
$z_{s}\leq\ell(\mu\boxplus\sigma_{s}).$
Because $\omega\leq\ell(\mu)$, $G_{\mu}(\omega)$ is non-positive and therefore $z_{s}$ is decreasing which shows that $z_{s}\geq z_{t}=H_{\mu,t}(B)=x$ for all $0\leq s\leq t$. This completes the proof of (A.6).
@@
Appendix B A priori properties on the rate function
@@
We may and shall assume that $\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-x|\leq\delta)$ is finite since otherwise we are done. The proof follows the proof of [CDG23, Lemma 5.2] based on [CDG23, Lemma 8.1] where a coupling is constructed. Because we restrict ourselves to Gaussian entries in this article, we propose a simpler version based on the symmetric Ornstein-Uhlenbeck
${\mathrm{d}G_{N}(s)=\mathrm{d}H_{N}(s)-\frac{1}{t}G_{N}(s)\mathrm{d}s}$
[…]
${\mathrm{d}(G_{N}(s)-\bar{G}_{N}(s))^{2}=-2(G_{N}(s)-\bar{G}_{N}(s))^{2}\mathrm{d}t}$
@@
$\|G_{N}(s)-\bar{G}_{N}(s)\|_{\infty}\leq\|G_{N}\|_{\infty}e^{-s}$ (B.1)
where $\|.\|_{\infty}$ denotes the operator norm. We will later take $s=N$. Moreover we work on the set where $\|G_{N}\|_{\infty}\leq N$ so that $\|G_{N}(t)-\bar{G}_{N}(t)\|_{\infty}$ is of order $Ne^{-N}\leq\delta/2$ if $t$ is of order $N$. We finally set $X_{N}(s):=B_{N}+G_{N}(s)$ and $\bar{X}_{N}(s):=B_{N}+\bar{G}_{N}(s)$.
[…]
$\mathbb{P}(|\mathbb{E}[\lambda_{1}(\bar{X}_{N}(N))]-\lambda_{1}(\bar{X}_{N}(N))|\geq\eta)\leq 2e^{-\frac{N\eta^{2}}{4}}{.}$ (B.2)
@@
Indeed, by (B.1) which implies that $\|X_{N}(N)-\bar{X}_{N}(N)\|_{\infty}$ is of order $Ne^{-N}$, for $N$ large enough
@@
where we used that $\bar{X}_{N}$ is independent from $X_{N}(0)=G_{N}$. Finally, we have seen that the above right-hand side goes to $1$ as $N$ goes to infinity. The fact that $Q_{N}$ is bounded from below by $1/2$ implies that
@@
Since $X_{N}(\frac{p+1}{N})-X_{N}(\frac{p}{N})$ is a GOE matrix with variance $t$, we find by (B.2) that $\mathbb{P}(\Omega_{N}^{c})$ is bounded above by $2N^{2}e^{-c(\delta)N^{2}}$ for some $c(\delta)>0$ provided $\delta>0$ is independent of $N$. Moreover, since $y$ lies in between $x$ and ${\ell}_{\nu,t}^{\Lambda}$, we see that
$\{|\lambda_{1}(X_{N}(N))-{\ell}_{\nu,t}^{\Lambda}|\leq\delta/2\}\cap\{|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N\}\cap\Omega_{N}\\ \subset{\bigcup_{1\leq p\leq N^{2}}\Bigr{\{}|\lambda_{1}(X_{N}(\frac{p}{N}))-y|\leq\delta\Bigr{\}}.}$
@@
$\begin{split}\mathbb{P}(|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N)&\leq 2\mathbb{P}(\cup_{1\leq p\leq N^{2}}\{|\lambda_{1}(X_{N}(\frac{p}{N}))-y|\leq\delta\})+4N^{2}e^{-c(\delta)N^{2}}\\ &\leq 2N^{2}\mathbb{P}(\{|\lambda_{1}(X_{N}(0))-y|\leq\delta\})+4N^{2}e^{-c(\delta)N^{2}},\end{split}$
@@
The first point is a slight generalization of the previous proof where we take the supremum over vectors $\mathsf{M}\in\mathcal{B}^{N}_{\delta}(\alpha)$. Moreover, because $G_{N}\mapsto\lambda_{1}(B_{N}+G_{N})$ is Lipschitz with Lispchitz constant $1$, we deduce as in [AGZ10, Theorem 2.3.5] that since $G_{N}$ has Gaussian entries with variance $1/N$, for every $\delta>0$,
$\mathbb{P}(|\lambda_{1}(B_{N}+G_{N})-\mathbb{E}[\lambda_{1}(B_{N}+G_{N})]|\geq\varepsilon)\leq 2e^{-\frac{\varepsilon^{2}}{4}N}\,.$

### v2

@@ 3.1. Properties of the rate function
We first prove the approximation and convexity properties of the rate function $I_{\nu,t}^{\Lambda}$ defined in (1.9). Recall that $\eta_{\nu}$ is the transport map such that $\nu=\eta_{\nu}\#1_{[0,1]}\mathrm{d}x$.
@@
Indeed the term in the r.h.s is clearly positive since
@@
$-\inf_{[x,+\infty)}I_{\nu_{\varepsilon}^{-},t}^{\Lambda}\leq\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\qquad\qquad\\ \qquad\qquad\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\leq-\inf_{[x,+\infty)}I_{\nu_{\varepsilon}^{+},t}^{\Lambda}.$
@@
$-\inf_{(x,+\infty)}I_{\nu,t}^{\Lambda}\leq\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\qquad\qquad\\ \qquad\qquad\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\geq x\right)\leq-\inf_{[x,+\infty)}I_{\nu,t}^{\Lambda}.$ (3.4)
@@
$-\inf_{(-\infty,x]}I_{\nu_{\epsilon}^{+},\Lambda}\leq\liminf_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\leq x\right)\qquad\\ \qquad\qquad\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}\left(\lambda_{1}(B_{N}+G_{N})\leq x\right)\leq-\inf_{(-\infty,x]}I_{\nu_{\varepsilon}^{-},\Lambda}.$
@@ 4.1. The case of the GOE with an outlier
$H_{0,t}(x)=x+\frac{t}{x}.$
Let $\omega_{0,t}$ be the inverse of $H_{0,t}$ on $(-\infty,-2\sqrt{t}]$, namely
@@
With the normalization of [Mai07], taking $\beta=1$ gives a LDP for the smallest eigenvalue of $G_{N}+\Lambda e_{1}e_{1}^{T}$ where $G_{N}$ is GOE of variance $t=1/2$. In the sequel we take $t=1/2$. Denote $\rho_{0,1/2}^{\Lambda}:=\Lambda+\frac{1}{2\Lambda}$. Let $J_{\Lambda}$ be the rate function found in [Mai07], given if $\Lambda\leq-\frac{1}{\sqrt{2}}$ by
@@
${M_{\Lambda}}(x):=-\frac{1}{2}\int_{x}^{\rho_{0,1/2}^{\Lambda}}\sqrt{z^{2}-2}\mathrm{d}z-\Lambda(x-\rho_{0,1/2}^{\Lambda})+\frac{1}{4}(x^{2}-\rho_{0,1/2}^{\Lambda})$
@@
$J_{\Lambda}(x):=\begin{cases}N_{\Lambda}(x)&\text{if $x\leq\rho_{0,1/2}^{\Lambda}$}\\ \int_{x}^{-\sqrt{2}}\sqrt{z^{2}-2}\mathrm{d}z&\text{if $\rho_{0,1/2}^{\Lambda}\leq x\leq-\sqrt{2}$}\\ +\infty&\text{if $x>-\sqrt{2}$},\end{cases}$
@@
${{N_{\Lambda}}(x):=\frac{1}{2}\int_{x}^{-\sqrt{2}}\sqrt{z^{2}-2}\mathrm{d}z-\Lambda x+\frac{1}{4}x^{2}+\frac{1}{4}+\frac{1}{4}\log 2+\frac{\Lambda^{2}}{2}.}$
@@
Let $x<-\sqrt{2}$. Assume that $\Lambda\leq-\frac{1}{\sqrt{2}}$, or $\Lambda\geq-\frac{1}{\sqrt{2}}$ and $x<-\sqrt{2}$. One can note that when $x\leq\rho_{0,1/2}^{\Lambda}$,
@@
Moreover, since $w_{0,1/2}=\gamma_{1}(x)$,
$\partial_{x}I_{0,1/2}^{\Lambda}(x)=\omega_{0,1/2}(x)-\Lambda=-\frac{1}{2}\sqrt{x^{2}-2}+\frac{x}{2}-\Lambda=J_{\Lambda}^{\prime}(x).$
Assume that $\Lambda\geq-\frac{1}{\sqrt{2}}$ and $x>\rho_{0,1/2}^{\Lambda}$. By Equation (3.2),
$\partial_{x}I_{0,1/2}^{\Lambda}(x)=\omega_{0,1/2}(x)-\omega_{0,1/2}^{*}(x)=-\sqrt{x^{2}-2}=J_{\Lambda}^{\prime}(x).$
@@ 4.2. The no outlier case
Let $\rho_{\nu,{1}}^{\Lambda}:=\ell_{\nu}+G_{\nu}(\ell_{\nu})\in\mathbb{R}\cup\{-\infty\}$. For all $x\leq\ell_{\nu,1}$, let $\gamma_{1}(x)\leq\gamma_{2}(x)$ be the two solutions of the equation $H_{\nu,1}(\gamma)=x$ on $(-\infty,\ell_{\nu})$. First observe that for all $x<\ell_{\nu,1}$,
@@ Appendix A Reminder on the free convolution
We recall some classical results on free convolution from [Bia97] that are useful throughout this article. Recall the definitions of $v_{t}(u)$ and $\Omega_{\mu,t}$ in (1.6) and (1.5). We have the following:
@@
(a) Characteristics and left edge of $\nu\boxplus\sigma_{t}$.
(b) The BBP transition.
In Figure A, gray lines represent characteristics $\{(H_{\nu,t}(x)=x+tG_{\nu}(x),t):t>0\}$ for given $x$ (given by the value on the horizontal axis). Characteristics starting around $(\omega_{\nu,t}(\ell_{\nu,t}),t)$ merge at $t$, see the discussion in Remark A.1, since this point corresponding to the merging of the two solutions. The blue curve is the set of merging points and gives the left edge $\ell_{\nu,t}$ of $\nu\boxplus\sigma_{t}$. Figure B illustrates the BBP transition. The $x$-coordinate of the gray dots represent the asymptotic smallest eigenvalue $\ell_{\nu,t}^{\Lambda}$ of $G_{N}+D_{N}$ where $G_{N}$ is a GOE of variance $t$ and $D_{N}$ as above, see Theorem 1.1. The BBP transition corresponds to the case when the gray dotes do not belong to the blue curve do not coincide.
@@
    Moreover, for $\eta_{1}<\cdots,\eta_{p}$, denoting $\nu_{\alpha}=\sum\alpha_{i}\delta_{\eta_{i}}$, we have for every $t>0$ and $\alpha,\alpha^{\prime}\in\Delta^{p}$,
@@
Assume first that $x\leq\ell_{\mu\boxplus\sigma_{t}}$. First by [GM20, Lemma 11], we have
$\omega_{\mu,t}((-\infty,\ell_{\mu\boxplus\sigma_{t}}])\subset(-\infty,\ell_{\mu}]$
and therefore $\omega\leq\ell_{\mu}$. This proves that $z_{0}\leq\ell_{\mu}$. Let $s\in(0,t)$. Recall that $\mu\boxplus\sigma_{t}=(\mu\boxplus\sigma_{s})\boxplus\sigma_{t-s}$. We have
@@
$z_{s}\leq\ell_{\mu\boxplus\sigma_{s}}.$
Because $\omega\leq\ell_{\mu}$, $G_{\mu}(\omega)$ is non-positive and therefore $z_{s}$ is decreasing which shows that $z_{s}\geq z_{t}=H_{\mu,t}(B)=x$ for all $0\leq s\leq t$. This completes the proof of (A.6).
@@
Appendix B A priori properties of the rate function
@@
We may and shall assume that $\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-x|\leq\delta)$ is finite since otherwise we are done. The proof follows the arguments of [CDG23, Lemma 5.2] based on [CDG23, Lemma 8.1] where a coupling is constructed. Because we restrict ourselves to Gaussian entries in this article, we propose a simpler version based on the symmetric Ornstein-Uhlenbeck
${\mathrm{d}G_{N}(s)=\mathrm{d}H_{N}(s)-\frac{1}{2t}G_{N}(s)\mathrm{d}s}$
where $H^{N}$ is a $N\times N$ symmetric Brownian motion and $G^{N}_{0}=G_{N}$ is a GOE with variance $t$. We also denote by $\bar{G}_{N}(t)$ the solution constructed with the same Brownian motion but starting from zero. It is well known that $G_{N}(s)$ has the law of a GOE with variance $t$ at any time $s$, whereas $\bar{G}_{N}(s)$ has the law of a GOE with variance $t(1-e^{-s/t})$. Moreover,
${\mathrm{d}(G_{N}(s)-\bar{G}_{N}(s))^{2}=-\frac{1}{t}(G_{N}(s)-\bar{G}_{N}(s))^{2}\mathrm{d}t}$
@@
$\|G_{N}(s)-\bar{G}_{N}(s)\|_{\infty}\leq\|G_{N}\|_{\infty}e^{-s/2t}$ (B.1)
where $\|.\|_{\infty}$ denotes the operator norm. We will later take $s=N$. Moreover we work on the set where $\|G_{N}\|_{\infty}\leq N$ so that $\|G_{N}(s)-\bar{G}_{N}(s)\|_{\infty}$ is of order $Ne^{-s/2t}\leq\delta/2$ if $s$ is of order $N$(here $t$ is fixed, independently of $N$). We finally set $X_{N}(s):=B_{N}+G_{N}(s)$ and $\bar{X}_{N}(s):=B_{N}+\bar{G}_{N}(s)$. Note that for every integer $p$, $X_{N}(\frac{p}{N})$ has the same law as $X_{N}(0)$ and in particular its smallest eigenvalue converges towards $\ell_{\nu,t}$ almost surely and in expectation. Observe also that by (B.1), $\|\bar{X}_{N}(N)-X_{N}(N)\|_{\infty}$ is bounded by $\|G_{N}\|_{\infty}e^{-N/2t}$ so that
$|\mathbb{E}[\lambda_{1}(\bar{X}_{N}(N))]-\mathbb{E}[\lambda_{1}(X_{N}(N))]|\leq\mathbb{E}[\|G_{N}\|_{\infty}]e^{-N/2t}$
goes to zero as $N$ goes to infinity. Hence, $\mathbb{E}[\lambda_{1}(\bar{X}_{N}(N))]$ converges as $N$ goes to infinity towards ${\ell}_{\nu,t}^{\Lambda}$. By concentration of measure for $\lambda_{1}(\bar{X}_{N}(N))$ (see [AGZ10, Theorem 2.3.5]), we also have that for every $\eta>0$,
$\mathbb{P}(|\mathbb{E}[\lambda_{1}(\bar{X}_{N}(N))]-\lambda_{1}(\bar{X}_{N}(N))|\geq\eta)\leq 2e^{-\frac{N\eta^{2}}{4t}}{.}$ (B.2)
@@
Indeed, (B.1) implies that $\|X_{N}(N)-\bar{X}_{N}(N)\|_{\infty}$ is of order $Ne^{-N/2t}$ when the operator norm of $G_{N}$ is bounded by $N$, for $N$ large enough so that $Ne^{-N/2t}\leq\delta/2$,
@@
where we finally used that $\bar{X}_{N}$ is independent from $X_{N}(0)=G_{N}$. Finally, we have seen that the above right-hand side goes to $1$ as $N$ goes to infinity. The fact that $Q_{N}$ is bounded from below by $1/2$ implies that
@@
Since $X_{N}(\frac{p+1}{N})-X_{N}(\frac{p}{N})$ is a GOE matrix with variance $t/N$, we find by (B.2) that $\mathbb{P}(\Omega_{N}^{c})$ is bounded above by $2N^{2}e^{-\frac{1}{4t}N^{2}\delta^{2}}$. Moreover, since $y$ lies in between $x$ and ${\ell}_{\nu,t}^{\Lambda}$, we see that
$\{|\lambda_{1}(X_{N}(N))-{\ell}_{\nu,t}^{\Lambda}|\leq\delta/2\}\cap\{|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N\}\cap\Omega_{N}\\ \subset{\bigcup_{1\leq p\leq N^{2}}\Bigr{\{}\Bigr{|}\lambda_{1}(X_{N}(\frac{p}{N}))-y\Bigr{|}\leq\delta\Bigr{\}}.}$
@@
$\begin{split}\mathbb{P}(|\lambda_{1}(X_{N}(0))-x|\leq\delta,\|G_{N}\|_{\infty}\leq N)&\leq 2\mathbb{P}(\cup_{1\leq p\leq N^{2}}\{|\lambda_{1}(X_{N}(\frac{p}{N}))-y|\leq\delta\})+4N^{2}e^{-c(\delta)N^{2}}\\ &\leq 2N^{2}\mathbb{P}(\{|\lambda_{1}(X_{N}(0))-y|\leq\delta\})+4N^{2}e^{-\frac{1}{4t}N^{2}\delta^{2}},\end{split}$
@@
The first point is a slight generalization of the previous proof where we take the supremum over vectors $\mathsf{M}\in\mathcal{B}^{N}_{\delta}(\alpha)$. Moreover, because $G_{N}\mapsto\lambda_{1}(B_{N}+G_{N})$ is Lipschitz with Lispchitz constant $1$, we deduce as in [AGZ10, Theorem 2.3.5] that since $G_{N}$ has Gaussian entries with variance $t/N$, for every $\delta>0$,
$\mathbb{P}(|\lambda_{1}(B_{N}+G_{N})-\mathbb{E}[\lambda_{1}(B_{N}+G_{N})]|\geq\varepsilon)\leq 2e^{-\frac{\varepsilon^{2}}{4t}N}\,.$